OpenStudy (anonymous):

Write an equation in slope-intercept form for the line passing through (–1, 10) and (5, 8).

OpenStudy (***[isuru]***):

do u know how to find the slope ?

OpenStudy (***[isuru]***):

\[slope = \frac{ y_{2} - y_{1} }{ x_{2} - x_{1}}\]

OpenStudy (anonymous):

is the slope -1/3?

OpenStudy (***[isuru]***):

\[slope = \frac{ 8 -10 }{ 5 - (-1) } = \frac{ -2 }{ 6 } = \frac{ -1 }{ 3 }\]

OpenStudy (anonymous):

okay. how do i find the Y intercept?

OpenStudy (***[isuru]***):

now the slope intercept form is given by y= mx + b where m is the slope and b is the intercept so.. since u now know the slope u can write it as y = (-1/3)x + b now to find b u can use the pair of coordinates ( 5 , 8) in (5 , 8) x = 5 and y = 8 if u plug them in to the equation y= (-1/3)x + b 8 = (-1/3)5 + b b = 29/3 so... the equation of the line in slope intercept form is y = (-1/3)x + 29/3 got it ;) ?

OpenStudy (anonymous):

yes! thank you sooo much :)

OpenStudy (***[isuru]***):

u r welcome!